OpenStudy (erniellyd28):

a rectangular is 42 feet longer than it is wide. its perimeter is 276 fee. (a) let w. be the width of the garden and L. be the length. write a system of linear equations whose solution gives the length and width of the garden. (b) use the method of substitution to solve the system.

OpenStudy (cruffo):

Are you interested in the way to solve the problem or just the answer?

OpenStudy (erniellyd28):

both please if possible so i don't have to keep bugging people with the same problem over and over lol

OpenStudy (cruffo):

Great Attitude! Since you are asked about a "configuration" (that is, problems involving geometric shapes) you may try starting by drawing the shape and thinking about what common things you know about that shape.

OpenStudy (cruffo):

For example: 

OpenStudy (cruffo):

Then look at the information the problem give you: "a rectangular is 42 feet longer than it is wide. its perimeter is 276 feet. " How can you use this information? 1.) How do you interpret the phrase " 42 feet longer than it is wide" in terms of L and W?

OpenStudy (erniellyd28):

2l+2w=276?

OpenStudy (cruffo):

Yes!!! That's the interpretation of the second piece of information!

OpenStudy (cruffo):

The first sentence: 42 feet longer than it is wide This can be one of two things: L = W + 42 or W = L + 42

OpenStudy (erniellyd28):

so either one would be correct?

OpenStudy (cruffo):

Excellent! In most problems like these the math will work itself out.

OpenStudy (cruffo):

So you now have two equations for your system: (part a done) 2L + 2W = 276 L = W + 42

OpenStudy (erniellyd28):

but do i get to the solution?

OpenStudy (cruffo):

That's the second part of your question. Using "substitution".

OpenStudy (cruffo):

See how the second equation is already solved for L...?

OpenStudy (erniellyd28):

yes

OpenStudy (cruffo):

You can "substitute" (W + 42) in for "L" in the first equation to get: 2(W + 42) + 2W = 276 then solve for W.

OpenStudy (erniellyd28):

this is where i get lost and or confused

OpenStudy (cruffo):

No problem. Do you recall a technique called "distribution" as a way of "getting rid of parentheses?

OpenStudy (erniellyd28):

yes

OpenStudy (cruffo):

To solve, start by distributing: 2(W + 42) + 2W = 276 2W + 84 + 2W = 276 notice: after distribution I don't include the ( )

OpenStudy (cruffo):

does that make sense ...?

OpenStudy (erniellyd28):

where did 84 come from?

OpenStudy (erniellyd28):

yes it is starting to make sense

OpenStudy (cruffo):

good question. 84 is 2 times 42. Distribute is just fancy for multiplying. 2(W + 42) ---> 2 * W + 2 * 42 ---> 2W + 84

OpenStudy (erniellyd28):

got it ;-)

OpenStudy (cruffo):

Great!

OpenStudy (cruffo):

So... 2W + 84 + 2W = 276 What do you do now?

OpenStudy (erniellyd28):

divide 2w on both sides?

OpenStudy (cruffo):

You could.... but that would get messy... Remember, when you divide both sides by anything you have to divide everything!! \[\frac{ 2W + 84 + 2W }{ 2W } = \frac{ 276 }{ 2W}\] which is the same as \[\frac{ 2W}{2W} + \frac{84}{2W} + \frac{2W }{ 2W } = \frac{ 276 }{ 2W}\] Hint: htere is a better way to proceed.

OpenStudy (cruffo):

May I suggest "gathering like terms" :)

OpenStudy (erniellyd28):

that sounds alot easier :-)

OpenStudy (cruffo):

I'm glad you think so! Usually, you should save division as the last step. So... 2W + 84 + 2W = 276 (combine the W's) 4W + 84 = 276

OpenStudy (erniellyd28):

there are so many steps to this

OpenStudy (cruffo):

It may seem that way, but we have only really taken three steps ... 2(W + 42) + 2W = 276 (first step was substitution) 2W + 84 + 2W = 276 (second step was distribution) 4W + 84 = 276 (third step was combining like term) (there are three more to finish this problem) You do much more complicated things everyday - for example, using a computer to talk about math online !)

OpenStudy (cruffo):

And when you work a bunch of these problems, the steps start making more sense so you don't second guess what step comes next.

OpenStudy (erniellyd28):

true that. so what is the next step?

OpenStudy (cruffo):

Let's get the term with the W in it by itself, that is, you want to have the following: 4W = (a number) but you have 4W + 84 = 276. How do you "move" or "cancel" the +84 ?

OpenStudy (erniellyd28):

divide 4w on both sides?

OpenStudy (cruffo):

If we had 4W = #, then dividing by 4 would be a great step. But we are not there yet,,,,

OpenStudy (cruffo):

What happens if we subtract 84 from both sides: 4W + 84 - 84 = 276 - 84

OpenStudy (erniellyd28):

276-84=192

OpenStudy (cruffo):

good. And what about the left side?

OpenStudy (cruffo):

4W + 84 - 84 would simplify how?

OpenStudy (erniellyd28):

4w+0?

OpenStudy (cruffo):

Yes... Is that the same as 4W (without the +0)

OpenStudy (erniellyd28):

yes?

OpenStudy (cruffo):

Yes!

OpenStudy (cruffo):

Example: 2 + 0 is just 2.

OpenStudy (erniellyd28):

yes

OpenStudy (cruffo):

Almost there.... 2(W + 42) + 2W = 276 (first step was substitution) 2W + 84 + 2W = 276 (second step was distribution) 4W + 84 = 276 (third step was combining like term) 4W = 192 (4th step was to subtract 84 from both sides) and the final step to solve for W is ... your favorite!!! divide both sides of the equation by 4

OpenStudy (erniellyd28):

lol

OpenStudy (erniellyd28):

i object to that remark lol

OpenStudy (erniellyd28):

jk

OpenStudy (cruffo):

:)

OpenStudy (erniellyd28):

192 divided 4 is 48

OpenStudy (cruffo):

Lovely! So now you know the width is 48. Look back on the original problem and the equations you created in part 1 of this problem. How can you find the length?

OpenStudy (erniellyd28):

divide 48 into 276?

OpenStudy (cruffo):

You are kinda obsessed with dividing things!

OpenStudy (erniellyd28):

sorry

OpenStudy (cruffo):

The original problem states that the length is 42 feet longer that the width.

OpenStudy (erniellyd28):

thats the only way i have been used to doing things aside from multiplying things

OpenStudy (cruffo):

We'll get to that in a sec...

OpenStudy (cruffo):

If you know that the width is 48 feet, and that the length is 42 feet longer than that.. then you can say that the length is .... 48 feet + 42 feet ...

OpenStudy (erniellyd28):

90?

OpenStudy (cruffo):

Why the ? You got it!

OpenStudy (erniellyd28):

no i wasnt sure if i was supposed to actually add them together

OpenStudy (cruffo):

sure, so long as the numbers have the same units. For example, If you have 2 apples and you get another 5 apples, then you have 7 apples total. 2 apples + 5 apples = 7 apples

OpenStudy (erniellyd28):

you are making me hungry lol.jk

OpenStudy (cruffo):

I know, right!

OpenStudy (cruffo):

Look... What you just had to solve was called a "multi-step linear equation" (I think you can guess why it's a multi-step). I would strongly suggest that you practice some "one-step" and "two-step" problems to make sure you master this "dividing both sides" "subtract from both sides" types of thing before continuing with problems like this. You would feel less confused, and more confident. Take a look at Khan Academy: videos and practice https://www.khanacademy.org/math/algebra-basics/core-algebra-linear-equations-inequalities

OpenStudy (erniellyd28):

i definitely need practice

OpenStudy (erniellyd28):

so whats the final step to this looooong process?

OpenStudy (cruffo):

Just to state the answer in terms of the problem: The rectangle is 48 feet by 90 feet.

OpenStudy (erniellyd28):

that would be the solution?

OpenStudy (cruffo):

Well there are two parts to the question: 1.) Create the system of equations answer was the two equations: 2L + 2W = 276 L = W + 42 2.) Use substitution to solve This was the long part: 2(W + 42) + 2W = 276 (first step was substitution) 2W + 84 + 2W = 276 (second step was distribution) 4W + 84 = 276 (third step was combining like term) 4W = 192 (4th step was to subtract 84 from both sides) W = 48 (5th step was to divide both sides by 4) Since W = 48, then substituting 48 in for W in the 2nd equation from part one you get: L = W + 42 L = 48 + 42 L = 90 So the final result is that the rectangle is 48 feet wide and 90 feet long.